OpenStudy (anonymous):

Evaluate the expression √(-4)√(-49) and write in the form of a+bi.

Directrix (directrix):

@swilson9 √(-4)√(-49) = i * √(4) * i * √(49) = i * 2 * I * 7 = ? Please simply the left side and post what you get. Someone will be by to check and help you get the result into a + bi form if you need it.